OpenStudy (aravindg):

Please help : Let X={1,2,3,....10}.Find the number of pairs {A,B} such that A and B are a subsets of X .But A not equal to B . A intersection B={5,6,8}.

OpenStudy (aravindg):

@Chlorophyll , @nincompoop, @UnkleRhaukus , @satellite73 , @.Sam. , @ajprincess , @apoorvk , @ash2326 , @asnaseer , @cwrw238 , @Directrix , @dumbcow , @eliassaab , @experimentX , @ganeshie8 , @hartnn , @Hero , @hba, @Ishaan94

OpenStudy (amistre64):

What do we know about set A?

OpenStudy (amistre64):

A and B are elements of X AnB share the elements 5,6,8 ok, i see it better now

OpenStudy (amistre64):

A={5,6,8} B={1,2,3,4,5,6,7,8,9,10} seems to be a set we can use; i wonder does the arrangement of elements of X between A and B make a difference? A={1,2,3,4,5,6,8} B={5,6,7,8,9,10} would be another suitable set AxB would seem to be a way to count all the possible ordered pairs to me 5 * * * * x * * * * * 6 * * * * * x * * * * 8 * * * * * * * x * * 1 2 3 4 5 6 7 8 9 10 3*10 - 3 = 27 for the first setup; now to see if the second setup makes a difference 10 9 8 x 7 6 x 5 x 1 2 3 4 5 6 8 it seems like there are always going to be 3 points of the set that are against the rules 6*7-3 = 39 ordered pairs so i spose arrangement of sets does make a difference

OpenStudy (anonymous):

how about A x B

OpenStudy (amistre64):

i believe the sets that create the most points is the 7x6 3*10 = 10 , 27 4*9 = 36 , 33 5*8 = 40 , 37 6*7 = 42 , 39 the number of ordered pairs that match the criteria are then

OpenStudy (amistre64):

3*10 = 30 :)

OpenStudy (anonymous):

we have two generic possibilities A = * * * * 5 6 * 8 * * B = * * * * 5 6 * 8 * * , where you can fill in numbers where * is

OpenStudy (anonymous):

might as well keep them in order, so the cases are , if you put a number in for * in A, then you must leave it out for B, and vice a versa

OpenStudy (anonymous):

, so lets say , for instance, put in 1 . we have two cases (A,B) A = 1 * * * * 5 6 * 8 * * A = * * * * 5 6 * 8 * * B = * * * * 5 6 * 8 * * B = 1 * * * 5 6 * 8 * *

OpenStudy (anonymous):

, so now you can proceed by singletons , 1,2,3.. then pairs, then triples

OpenStudy (shubhamsrg):

RMO 2012 hmm,,did you give? i have seen the question paper..

OpenStudy (anonymous):

actually theres more cases thatn i anticipated

OpenStudy (amistre64):

are we to find the total number of ordered pairs, as in all the possible sets that can be constructed? or is it the most ordered pairs from a maximal set?

OpenStudy (anonymous):

ok thats the best way, count by |A|= 3, |B|=3 , then |A| = 4, |B|=3 , etc

OpenStudy (anonymous):

, |A| means cardinality of A

OpenStudy (anonymous):

for |A|= 3, |B|=3, there is just one case for (A,B)

OpenStudy (anonymous):

now do |A| = 3, |B| = 4, how many (A,B) pairs are there

OpenStudy (anonymous):

there are 7 pairs when |A| = 3 , | B| = 4 ,

OpenStudy (shubhamsrg):

A and B will surely have 5,6,8 we are left with 1,2,3,4,7,9,10 1st case.. 0 elements in A or B for 0 elements in A, B can have 1,2,3..or 7 elements.. =>(7C1 + 7C2 ... 7C7) we multiply this by 2 since A and B are symmetrical i.e. same thing will take place for 0 elements in B so for 1st case : (7C1 +7C2... 7C7)*2 case 2// 1 element in A or B =>(6C1 + 6C2..6C6)* 7 *2 case 3.. 2 elements in A or B 7C2 * (5C2 + 5C3 .. 5C5) * 2 case 4: 3 elements in A or B 7C3 * (4C3 + 4C4) *2 case 5: wont be there.. i havent elaborated my solution since i have an intuition it may be wrong.. my ans comes out to be sum of all cases..anyone please check mine..

OpenStudy (anonymous):

ok here is an exhaustive list, a schemata

OpenStudy (anonymous):

Let |A| = m |B| = k m= 3 k = 3 m=3 k = 4 m=3 k = 5 ... m=3 k = 7 m = 4 k = 3 m= 4 k = 4 m=4 k = 5 . . . m = 4 k = 7 until you get to m= 7 k = 3... m = 7 k = 7

OpenStudy (anonymous):

hey do you hvae the final answer to check ?

OpenStudy (anonymous):

I got 7C0 ( 7C0 + 7C1 + 7C2 + 7C3 + 7C4) + 7C1 ( 7C0 + 7C1 + 7C2 + 7C3 + 7C4) 7C2 ( 7C0 + 7C1 + 7C2 + 7C3 + 7C4) + 7C3 ( 7C0 + 7C1 + 7C2 + 7C3 + 7C4) + 7C4 ( 7C0 + 7C1 + 7C2 + 7C3 + 7C4) = ( 7C0 + 7C1 + 7C2 + 7C3 + 7C4) ^2

OpenStudy (anonymous):

9801

OpenStudy (aravindg):

@eddie2b SRRY i dont have final answer to check

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you understand my explanation ?

OpenStudy (aravindg):

@amistre64 , @eddie2b landed on the answer !! its 2186!!

OpenStudy (aravindg):

i made a small editing in the question check:

OpenStudy (anonymous):

A is not equal to B , ok

OpenStudy (anonymous):

so the answer is 2186, are you sure?

OpenStudy (sirm3d):

there are \[\large 2^{10}=1024\]distinct subsets of X={1,2,...,10}

OpenStudy (aravindg):

yes!!

OpenStudy (sirm3d):

if we form pairs, then it is just a combination problem with n = 1024 and r = 2

OpenStudy (aravindg):

but i dont understand hw to get it !!

OpenStudy (anonymous):

are you sure its 2186

OpenStudy (sirm3d):

\[\large \left(\begin{matrix}1024 \\ 2\end{matrix}\right) =\frac{ 1024 \times 1023 }{ 2 \times 1 }=512 \times 1023\]

OpenStudy (aravindg):

yes i even got the solution but i dont understand it!!check question 4

OpenStudy (anonymous):

one moment

OpenStudy (aravindg):

thanks to @shubhamsrg i searched online for rmo 2012 and gt this bt the solution is nt clear

OpenStudy (anonymous):

use slots

OpenStudy (anonymous):

A = _ _ _ _ 5 _ 7 8 _ _ B = _ _ _ _ 5 _ 7 8 _ _

OpenStudy (aravindg):

?

OpenStudy (sirm3d):

i understand to solution. why bother about 5, 7 and 8 when they are in BOTH sets? just focus on whether to put the rest of the digits or not.

OpenStudy (aravindg):

@sirm3d can u xplain me nt getting it

OpenStudy (anonymous):

looks like a typo is in your solutio

OpenStudy (sirm3d):

consider the digit 1. will you put it in A or in B or not? that gives you three choices.

OpenStudy (anonymous):

rest of the elements can be assigned in *3* ways

OpenStudy (anonymous):

who wrote this solution?

OpenStudy (aravindg):

experts from resonance institute india

OpenStudy (anonymous):

aravind, ok start by picking an element , say 1

OpenStudy (sirm3d):

each of the seven numbers gives you 3 choices, or a total of 3^7=2187 choices.

OpenStudy (zarkon):

why cant 1 be in both A and B

OpenStudy (anonymous):

A = _ _ _ _ 5 _ 7 8 _ _ B = _ _ _ _ 5 _ 7 8 _ _ for number 1, there are three choices, it is either in A , or B , or neither

OpenStudy (shubhamsrg):

what is the ans ? i am getting 2130 ..

OpenStudy (sirm3d):

@Zarkon the intersection would then include the digit 1 if 1 is in both sets A and B.

OpenStudy (aravindg):

@shubham check the pdf i posted its 2186

OpenStudy (zarkon):

oh i see the intersection is {5,6,8}...guess I should read

OpenStudy (anonymous):

aravind, so start with A = {5,7,8} , B = {5,7,8} , now the number 1 has 3 options , it can go in set A, or it can go in set B , or it can go in neither .

OpenStudy (aravindg):

A NOT EQUAL TO B?

OpenStudy (anonymous):

so we are looking at the set { 1,2,3,4,6,9,10} ,

OpenStudy (aravindg):

wasnt that a ondition?

OpenStudy (anonymous):

there is only case where they will be equal , and so you subtract 1 from the final

OpenStudy (anonymous):

so the number 1 has three options where it can go , and then for each of that case the next case is number 2 has three options where it can go, and so on , multiply 3^7 , then subtract 1 for the case when you have none in both

OpenStudy (anonymous):

how about this, let's label the points in the set

OpenStudy (anonymous):